1. Suppose that each firm in a perfectly competitive market has a short-run total cost of TC = 75 + 500Q – 5Q2 + 0.5Q3, where MC = 500 – 10Q + 1.5Q2.


  1. Calculate the output that minimizes the firm’s AVC.
  2. What is the firm’s shutdown price?




1
Expert's answer
2021-03-04T07:24:51-0500

(a) Total cost is the sum of fixed cost and variable cost:


"TC=FC+VC."

In our case the variable cost of the firm takes form:


"VC=500Q-5Q^2+0.5Q^3."

Let's find average variable cost (AVC). By the definition,


"AVC=\dfrac{VC}{Q}=500-5Q+0.5Q^2."

Let's take the derivative of AVC equation:


"\dfrac{dAVC}{dQ}=-5+Q=0,""Q=5."

The output of 5 units minimizes the firm’s AVC.

(b) The shutdown price occurs at the minimum of the average variable cost curve at a point where MC=AVC:


"500-10Q+1.5Q^2=500-5Q+0.5Q^2,""Q^2-5Q=0,""Q(Q-5)=0,""Q=5."

In a perfectly competitive market P=MC, therefore:


"P=500-10Q+1.5Q^2."

Let's substitute "Q" into the previous equation and find the shutdown price:


"P=500-10\cdot5+1.5\cdot(5)^2=\$412.5"